# Vision-Language Models can Identify Distracted Driver Behavior from Naturalistic Videos

### v1

@@ V-C Hyperparameters and Ablation Studies
We present the performance of the Multi-frameCLIP model tested at various sampling rates on the SynDD1 dataset in Table X. We perform a 7-fold cross-validation experiment with different sampling rates on the Multi-frameCLIP model and other commonly used distracted driving backbones to see the effect of sampling rates on model performance. The results indicate that the CLIP-L/14-based Multi-frameCLIP achieves the highest top-1 accuracy at 1 FPS sampling rate, with an accuracy of 70.36%. This was the highest accuracy achieved across all tested models and sampling rate settings for this experiment. In comparison, the lowest accuracy values were consistently provided by AlexNet and Vgg16 models across all FPS settings, with both models achieving an accuracy of 12.62%. The results also suggest that varying the sampling rate can impact the performance of the models. For example, the performance of the CLIP-L/14 model decreased from 70.36% accuracy at 1 FPS to 62.08% at 5 FPS and to 61.25% at 20 FPS. These findings suggest that the optimal sampling rate may vary depending on the specific model employed. The changes from one frame to the next are more significant at a 1 FPS sampling rate, which helped the model make more accurate predictions. Therefore, 1 FPS is the optimal sampling rate for the Multi-frameCLIP model under the conditions of our study.
Table XI shows the performance of the VideoCLIP model on the SynDD1 dataset at different sampling rates of 1 FPS, 2 FPS, 5 FPS, and 20 FPS. We perform seven-fold cross-validation on the VideoCLIP model to see the effect of sampling rates on model performance. The ‘Mean’ column reports the average accuracy across all folds for each respective FPS setting. The results indicate that VideoCLIP achieves the highest Top-1 accuracy at 20 FPS, with an accuracy of 81.85%. A higher FPS results in a better temporal resolution, allowing the model to capture more subtle changes and movements that occur between frames. Consequently, the model may be better equipped to understand and predict based on these small changes. However, the exact impact of sampling rate would depend on the specific characteristics of the dataset and the downstream task.
@@
We illustrate the class-level precision, recall, and F1-scores of Single-frameCLIP on the SynDD1 dataset in Table XII. In this table, we analyze all eight classes to understand the most challenging actions to classify. In addition, we demonstrate the confusion matrix of the action classes in Fig. 8. We have found that classes such as “eating,” “singing and smiling,” and “yawning” exhibited lower performance metrics, indicating that these categories pose greater challenges for the framework to classify correctly. These action classes appeared to generate similar image features, leading to a higher rate of misclassification. To further illustrate this, Fig. 9 provides visual examples of instances where the Single-frameCLIP model incorrectly classified actions, comparing the model’s outputs against the ground truth.
TABLE XII: Single-frameCLIP: Precision, recall, and F1-score on SynDD1
Action classes Precision Recall F1-score
adjusting hair 0.51 0.74 0.60
drinking water 0.70 0.86 0.78
eating 0.60 0.31 0.41
picking up something 0.76 0.96 0.85
reaching behind 0.79 0.94 0.86
singing and smiling 0.44 0.30 0.36
talking to the phone 1.00 0.77 0.87
yawning 0.56 0.42 0.48
Figure 8: Confusion matrix (normalized) of Single-frameCLIP model on SynDD1 dataset. It shows the classes - “eating”,“yawning” and “singing” - for which Single-frameCLIP struggles to classify the distracted driving action.
TABLE XIII: Single-frame CLIP: Removing vs. merging classes results on SynDD1 dataset
@@
3)  Merging vs. Removing Classes
To fully understand the impact of the challenging classes on overall model performance, we pursue two additional experiments: merging and removing classes. Our initial focus was to explore how the overall model performance is influenced by either merging or removing classes that are more challenging to predict accurately. From the previous analysis, we identify that the Single-frameCLIP model has difficulty in accurately predicting the “singing”, “eating” and “yawning” classes compared to others. Table  XIII reflects our findings, which are categorized into three different experimental settings. The first row shows the outcome following the removal of the three challenging classes with three camera views. This modification resulted in an average F1 score of 0.763 and a mean accuracy of 75.62%. The second row illustrates the outcome of merging the three challenging classes with three camera views. This scenario led to a mean F1 score of 0.698 and an average accuracy of 69.41%. The third row provides the results obtained from an experiment in which the three challenging classes were removed and only the dashboard camera view was considered. This experimental setup led to an average F1 score of 0.873 and a mean accuracy of 87.08%. In conclusion, the optimization of the camera views, combined with the adjustment of the challenging classes, resulted in a significant improvement in the model’s performance.
@@
TABLE XIV: Single-frameCLIP: Effect of Camera-view on SynDD1 dataset
No. of class Camera-view Top-1 accuracy
5 Dashboard camera view 87.640
5 Rear mirror camera view 85.811
5 Right window camera view 76.056
In this section, we investigate the effect of camera position on our framework’s performance. In Table XIV, we present the Single-frameCLIP accuracy for five distracted classes, considering three distinct camera views. The outcomes of these experiments suggest that the use of a single camera view, particularly the dashboard view, yields superior results when compared to other camera views on the SynDD1 dataset. These results underscore the importance of camera perspective in driver distraction detection tasks. The dashboard camera view appears to provide the most effective angle for capturing driver behaviors due to the direct line of sight and better visibility of driver action.
V-D Anecdotal Observation
@@
We report some of the miss-classified frames by the Single-frameCLIP model on the SynDD1 dataset in Fig. 9 and the class-level confusion matrix in Fig 8. Analyzing the confusion matrix and the miss-classified frames, we conclude that the most challenging frames reside in the overlap of the three classes -“eating”, “yawning” and “singing” in the SynDD1 dataset. In addition, we manually look into the video files of the misclassified actions and noticed that these three classes are mostly dependent on specific hand or facial gestures. Most of the time, it is quite complicated to decipher these activities from each other by looking at only a single frame. To address this complexity, the Multi-frameCLIP and VideoCLIP frameworks with a majority voting scheme were developed, wherein a stack of frames is assessed for a more accurate prediction score. This consideration contributed to the noted performance enhancement relative to the Single-frameCLIP model.
Figure 9: Some example frames that were misclassified by the frameCLIP on SynDD1 dataset- (a) predicted: yawning, ground-truth: eating (b) actual yawning, (c) predicted: yawning, ground-truth: singing (d) actual yawning, (e) predicted: singing,ground-truth: yawning (f) actual singing
2)  Frame-by-frame Analysis for Video Input
@@
Figure 10: This plot shows the predicted probability score vs. frame number. We analyze the Multi-frameCLIP and VideoCLIP model’s prediction probability scores for the entire action. We notice distinct changes in frames for high and low probability scores. This figure demonstrates that the VideoCLIP outperforms Multi-frameCLIP’s prediction scores offering a better approach to capturing temporal features. Therefore, VideoCLIP performs better in classifying actions on video data
@@
Fig. 10 presents the frame-by-frame analysis of video input for driver ID 25470 from the SynDD1 dataset using the Multi-frameCLIP and VideoCLIP models. The depicted plot conveys the dynamic probabilities of predictions produced by the Multi-frameCLIP and VideoCLIP models over the entire span of action. The vertical axis shows the moving average of the computed probabilities in relation to the individual frame numbers, represented by the horizontal axis. We observe distinct fluctuations between high and low probability scores, aligning with frames where the model either successfully or unsuccessfully classified the given action. Our analysis focused on two distinct actions - adjusting hair and singing, at a sampling rate of 20 FPS. The representation of lower and higher probability scores from both models is clearly distinguished in this visualization. An important observation is the superiority of VideoCLIP’s prediction score over the entire course of action, compared to that of the Multi-frameCLIP model. Additionally, we notice that VideoCLIP demonstrates a stable and consistent performance, capturing low-probability frames and maintaining a consistent score throughout the video. In contrast, the Multi-frameCLIP model appears to struggle at certain frames. Therefore, we conclude that VideoCLIP has the capability to efficiently capture temporal information from videos significantly improving prediction accuracy while maintaining consistency. Moreover, the plot illustrates the superiority of both Multi-frameCLIP and VideoCLIP in extracting the temporal correlations among frames, thus providing insights into continuous actions that are challenging to get from isolated frames, as in the Single-frameCLIP scenario.
@@ VI Conclusion
In this paper, we have introduced a vision-language modeling framework aimed at categorizing distracted driving behaviors from naturalistic driving data. Performance of various CLIP-based frameworks have been evaluated on two distracted driving datasets, namely SynDD1 and StateFarm. Compared to the Zero-shotCLIP model, the Single-frameCLIP model benefited from task-based fine-tuning with significant improvement in accuracy while failing to capture the temporal dynamics of a video. Further experiments have shown that Multi-frameCLIP and VideoCLIP can leverage temporal information from video data, which allows for a better understanding of actions, events and sequences within a video. VideoCLIP and Multi-frameCLIP frameworks have outperformed Single-frameCLIP by a significant margin. Overall, our proposed frameworks exhibit state-of-the-art performance at distracted action classification, as tested on two public datasets, and outperform traditional and widely used models. Our findings have also shown that, unlike the other models assessed, which displayed a decrease in accuracy, the Single-frameCLIP model could maintain its performance levels consistently even when the quantity of training data was reduced to 20% of the original amount. This study has also evaluated the efficacy of different CLIP-model variants, highlighting that their performance can be influenced by a variety of factors including variable training dataset size, sampling rates, and camera perspectives. These observations can contribute to the development of more robust driver assistance systems to understand driver distractions. This study emphasizes the real-world applicability and generalizability of these models, while also pointing out areas for improvement, such as managing difficult classification categories and fine-tuning hyperparameters. Several future works could extend and refine our current CLIP-based driver distraction identification model. First, exploring the optimization of prompts for Vision-Language foundation models may enable the framework to extract more nuanced visual-textual relationships, leading to an improved classification. Second, the fusion of different camera views can be investigated as a means of providing a more comprehensive picture of driver behavior, potentially contributing to more accurate decision-making.
@@
Moreover, the generalizability of our model across various scenarios and datasets presents another intriguing prospect. For instance, assessing whether the model could be fine-tuned on one dataset and subsequently deployed on another could further illustrate its versatility and effectiveness. Finally, systematic evaluations under diverse environmental conditions, such as varying lighting situations and data quality, could offer valuable insights into the model’s robustness and ability to function optimally under real-world driving circumstances.

### v3

@@ V-C Hyperparameters and Ablation Studies
To see the effect of sampling rates on model performance, we test Multi-frameCLIP and VideoCLIP at various sampling rates on the SynDD1 dataset. Table XIII and Table XIV show the results. We maintain the same experimental settings and observe distinct trends across different sampling rates. The ViT-L/14-based Multi-frameCLIP achieves the peak Top-1 accuracy at a 1 FPS sampling rate with an accuracy of 70.36%. This was the highest accuracy achieved across all tested models and sampling rate settings for this experiment. In comparison, the AlexNet and VGG16 models got the lowest accuracy values. However, in Table XIV the VideoCLIP achieves the highest Top-1 accuracy at 20 FPS, with an accuracy of 81.85%. This contrast highlights the models’ sensitivity to sampling rate variations, suggesting that the optimal rate depends on the specific model and dataset characteristics. The changes from one frame to the next are more significant at a 1 FPS sampling rate, which helped the Multi-frameCLIP model make more accurate predictions. Therefore, 1 FPS is the optimal sampling rate for the Multi-frameCLIP model under the conditions of our study. On the other hand, as the average action duration for the SynDD1 dataset is $\sim$18.7 sec, a 20FPS rate results in a better temporal resolution, allowing the VideoCLIP model to capture more subtle changes and movements. If the sampling rate is too small, then the video clip will span more than the average action duration and confuse the model prediction. Therefore, the impact of the sampling rate would depend on the average action duration and the characteristics of the downstream tasks.
@@
We illustrate the class-level F1-scores of Single-frameCLIP on the DMD and SAM-DD datasets in Table XII. In this table, we analyze the ten classes of the datasets to understand the most challenging actions to classify. The average F1-scores on DMD and SAM-DD datasets are 0.819 and 0.847, respectively. In addition, we demonstrate the confusion matrix of Single-frameCLIP on fold02 of the DMD dataset as an example in Fig. 8. We have found that the DMD classes such as “reaching to backseat”, “reaching to side” and “driving safely” exhibited lower F1-scores ($<$0.70), indicating that these categories pose greater challenges for the framework to classify correctly. These action classes appeared to generate similar image features, leading to a higher rate of misclassification. Additionally, we noticed that only two SAM-DD classes “head dropping” and “touching hair” got F1-scores below 0.70 among the ten classes. To further illustrate this, Fig. 9 provides visual examples of instances where the Single-frameCLIP model incorrectly classified actions, comparing the model’s outputs against the ground truth.
Figure 8: Confusion matrix of Single-frameCLIP model on DMD dataset. It shows that the “reaching to backseat”, “reaching to side” and “driving safely” classes were the most challenging for the Single-frameCLIP model. Also, the dataset has limited “yawning” samples compared to the other classes. Additionally, we noticed that only two SAM-DD classes “head dropping” and “touching hair” got F1-scores below 0.70 among the ten classes.
TABLE XV: Single-frame CLIP: Removing vs. merging classes results on SynDD1 dataset
@@
TABLE XVI: Effect of Camera-view on SAM-DD dataset
Models Camera Top-1 Leaving a set
(ours) view accuracy of driver out
Single-frameCLIP Dashboard 88.919
Side-view 89.142
Multi-frameCLIP Dashboard 88.908
Side-view 89.142
3)  Merging and Removing Classes
To fully understand the impact of the challenging classes on overall model performance, we pursue two additional experiments: merging and removing classes. Our initial focus was to explore how the overall model performance is influenced by either merging or removing classes that are more challenging to predict accurately. From our previous analysis, we noticed that the Single-frameCLIP model has difficulty in accurately predicting the “singing”, “eating” and “yawning” classes on the SynDD1 dataset compared to other classes. Table XV reflects our findings, which are categorized into three different experimental settings. The first row shows the outcome following the removal of the three challenging classes with three camera views. This modification resulted in an average F1 score of 0.763 and a mean accuracy of 75.62%. The second row illustrates the outcome of merging the three challenging classes with three camera views. This scenario led to a mean F1 score of 0.698 and an average accuracy of 69.41%. The third row provides the results obtained from an experiment in which the three challenging classes were removed and only the dashboard camera view was considered. This experimental setup led to an average F1 score of 0.873 and a mean accuracy of 87.08%. In conclusion, the optimization of the camera views, combined with the adjustment of the challenging classes, resulted in a significant improvement in the model’s performance.
@@
In this section, we investigate the effect of camera position on our framework’s performance. In Table XVI, we present the Top-1 accuracy for Single-frameCLIP and Multi-frameCLIP on ten distracted classes for the SAM-DD dataset, considering two distinct camera views. The outcomes suggest that the side view yields superior results when compared to the dashboard camera view. These results underscore the significance of the camera’s perspective in detecting driver distractions. The side camera view appears to provide the most effective angle for the SAM-DD dataset, capturing the driver’s movements and body maneuvers.
V-D Computational Complexity
TABLE XVII: VideoCLIP computation efficiency (Batch size=1)
FLOPs #Params. Forw./backw. Params Estimated
(G) (M) pass size(MB) size (MB) Total size(MB)
3.54 9.67 964.26 38.68 1021
In Table XVII, we present the details of the computational efficacy of the proposed VideoCLIP model. In terms of driver distraction recognition tasks, we compared the floating point operations (FLOPs) and the number of parameters (##\##Params.) of VideoCLIP with its counterparts. VideoCLIP demonstrates superior FLOPs efficiency and a relatively smaller parameter count ($<$10M). For example, in the SAM-DD dataset [4] setting, Swin-T gets 4.5G FLOPs and 28.3M trainable parameters. In [69] setting, VideoMAE operates at 57G FLOPs with 22M parameters, ResNet50 [42] uses 3.8 GFLOPs with 25.6M parameters for an input size of 224x224x3. Therefore, the computational performance of the VideoCLIP model coupled with a Top-1 accuracy of 81.85% makes it a resource-effective model in the driver distraction recognition task context. This efficiency is particularly valuable in real-time Intelligent Transportation applications like driver distraction recognition, where quick and accurate video processing is essential for safety and effectiveness. Moreover, the small parameter count also implies potentially lower memory requirements and faster training times, further enhancing its applicability in practical scenarios.
V-E Anecdotal Observation
@@
We demonstrate some miss-classified frames from the Single-frameCLIP model on DMD and SynDD1 dataset in Fig. 9 and the class-level confusion matrix of DMD classes in Fig 8. Analyzing the confusion matrix and the miss-classified frames, it is evident that the most challenging classes for the model occur where there is considerable feature overlap between classes that exhibit lower F1-scores. The middle column of Fig 9 shows the overlapping cues between two actions characterized by specific hand movements or facial expressions.
Figure 9: Some example frames that were misclassified by the Single-frameCLIP on DMD (top row) and SynDD1 (bottom row). These examples show evidence of how Single-frameCLIP fails to capture the temporal information only by looking at a single frame at a time.
For example, the first-row middle frame driver’s left-hand position close to the steering wheel resembles the starting motion to the side, which confused the model to incorrectly predict the “reaching side” class. Therefore, the Single-frameCLIP model struggles to distinguish actions where the visual movements are subtle and there is a slight temporal context. Most of the time, it is quite complicated to decipher these activities only by looking at a single frame. To address this issue, the Multi-frameCLIP and VideoCLIP frameworks were developed, where a sequence of consecutive frames is assessed for a more accurate prediction score. Moreover, these models leverage the temporal information across multiple frames and capture subtle changes in hand movements and facial gestures. This consideration contributed to a significant performance enhancement relative to the Single-frameCLIP model.
2)  Frame-by-frame Analysis on SynDD1 Video Clips
@@
Figure 10: This plot shows the predicted probability score vs. frame number. We analyze the Multi-frameCLIP and VideoCLIP model’s prediction probability scores for the entire action. We notice distinct changes in frames for high and low probability scores. This figure demonstrates that VideoCLIP outperforms Multi-frameCLIP’s prediction scores, offering a better approach to capturing temporal features. Therefore, VideoCLIP performs better in classifying actions on video data
@@
Fig. 10 presents the frame-by-frame analysis of video input for driverID 25470 from the SynDD1 dataset using the Multi-frameCLIP and VideoCLIP models. The depicted plot conveys the dynamic probabilities of predictions produced by the Multi-frameCLIP and VideoCLIP models over the entire span of action. The vertical axis shows the moving average of the computed probabilities in relation to the individual frame numbers, represented by the horizontal axis. We observe distinct fluctuations between high and low probability scores, aligning with frames where the model either successfully or unsuccessfully classified the given action. Our analysis focused on two distinct actions - adjusting hair and singing, at a sampling rate of 20 FPS. The representation of lower and higher probability scores from both models is clearly distinguished in this visualization. An important observation is the superiority of VideoCLIP’s prediction score over the entire course of action compared to that of the Multi-frameCLIP model. Additionally, we notice that VideoCLIP demonstrates a stable and consistent performance, capturing low-probability frames and maintaining a consistent score throughout the video. In contrast, the Multi-frameCLIP model appears to struggle at certain frames. Therefore, we conclude that VideoCLIP has the capability to efficiently capture temporal information from videos, significantly improving prediction accuracy while maintaining consistency. Moreover, the plot illustrates the superiority of both Multi-frameCLIP and VideoCLIP in extracting the temporal correlations among frames, thus providing insights into continuous actions that are challenging to get from isolated frames, as in the Single-frameCLIP scenario.
V-F Limitaions and Future directions
Our algorithm shows promising results within its current setup, yet there is potential to further its adaptability and performance in varied real-world driving conditions across different unusual distraction categories. The variety of distractions that drivers encounter is vast and often unpredictable. Therefore, we need a sophisticated monitoring system to recognize the natural interactions between humans and vehicles and encourage safe driving practices. This system must be equipped to accurately identify the driver’s current state and assess the reliability of its predictions. Our current models were tested under controlled and available distracted settings from public datasets. The performance and efficacy in real-world applications, considering diverse hardware and environmental conditions, were not fully explored. To enhance the robustness and applicability of our algorithm, future work could involve training with additional data that cover a wider range of driving conditions - multiple distracted behaviors occurring simultaneously or alternately, unexpected or out-of-distribution behavior categories, occlusion events and so on. This would help in validating the algorithm’s effectiveness across various scenarios. Optimizing the algorithm for efficient performance across a range of hardware platforms, including low-power devices or with limited compute resources, could be a significant area of focus. This would make the system more viable for real-time applications in vehicles. Additionally, future research should also consider the ethical implications and privacy concerns associated with driver monitoring, ensuring compliance with relevant regulations and societal norms. Also, we plan to address the misclassification issues (discussed in V-E1) and enhance the reliability of the model predictions by implementing reliable uncertainty estimates [70]. This would be particularly beneficial in complex or ambiguous scenarios where driver behavior is not distinctly classifiable, thus allowing for a more nuanced understanding and interpretation of the model’s confidence measures. We believe that this addition will enhance the safety and reliability of the automated driver assistance systems by providing a more comprehensive and trustworthy analysis of driver behavior.
@@ VI Conclusion
In this paper, we have introduced a vision-language modeling framework aimed at categorizing distracted driving behaviors from naturalistic driving data. The performance of various CLIP-based frameworks has been evaluated on four common distracted driving datasets. Compared to the Zero-shotCLIP model, the Single-frameCLIP model benefited from task-based fine-tuning with significant improvement in accuracy while failing to capture the temporal dynamics of a video. Further experiments have shown that Multi-frameCLIP and VideoCLIP can leverage temporal information from video data, which allows for a better understanding of actions, events and sequences within a video. VideoCLIP and Multi-frameCLIP frameworks have outperformed Single-frameCLIP by a significant margin. Overall, our proposed frameworks exhibit state-of-the-art performance at distracted action classification and outperform traditional and most widely used models. This study has also evaluated the efficacy of different CLIP-model backbones, highlighting that a variety of factors, including variable training dataset size, sampling rates, and camera angles, can influence their performance. These observations can contribute to the development of more robust driver assistance systems to understand driver distractions. This study emphasizes the real-world applicability and generalizability of these models while also pointing out areas for improvement, such as handling unexpected and out-of-distribution driving behavior categories and fine-tuning hyperparameters. Several future works could extend and refine our current CLIP-based driver distraction recognition model. First, exploring the optimization of prompts for Vision-Language foundation models may enable the framework to extract more nuanced visual-textual relationships, leading to an improved classification. Second, the fusion of different camera views can be investigated to deal with occlusion, provide a more comprehensive picture of driver behavior and make accurate decision-making.
@@
Moreover, the generalizability of our model across various scenarios and datasets presents another intriguing prospect. For instance, assessing whether the model could be fine-tuned on one dataset and subsequently deployed on another could further illustrate its versatility and effectiveness. In addition, we recognize the importance of empirical validation and testing our system under more challenging, unexpected, and out-of-distribution driving behavior categories. To this end, we are actively planning an extension of our current work. This next phase will include a series of experiments specifically designed to evaluate our framework’s response to real-world out-of-distribution actions. We also aim to extend the model’s testing to include a broader spectrum of actions, particularly focusing on those that are rare or challenging to classify. These tests are intended to validate our framework’s robustness and enhance its adaptive capabilities further. Finally, systematic evaluations under diverse environmental conditions, such as varying lighting conditions and data quality, could offer valuable insights into the model’s robustness and ability to function optimally under real-world driving circumstances.